Suppose 12 out of 30 families on your street have a cat or a dog as a pet.
a. what is the experimental probability that a randomly selected family in your neighborhood will have a cat or a dog as a pet?
b. based on P(cat or dog) from part (a), predict how many cat- or dog-owning families you can expect amoung 57 families in your neighborhood.
1
Expert's answer
2010-12-05T19:04:50-0500
a. 12 out of 30 families have a pet (a cat or a dog), thus the probability is P = 12/30 = 2/5 = 0.4.
b. There are 57*0.4 = 22.8, 22 families should be expected to have a pet.